# Generative Topological Networks

### v1

@@ 3 Experiments
In light of the discussion in Section 2.1 on using the intrinsic dimension of the data (or lower), we use autoencoders to represent each dataset in a lower dimensional latent space before training $\hat{h}$. Each dataset was trained separately (with its own autoencoder and $\hat{h}$). In all experiments, we use the same autoencoder architecture – a simple vanilla autoencoder (the encoder consisted of $2$ convolution layers with ReLU activation followed by a fully connected layer). For $\hat{h}$ we use a simple feedforward neural network with the width and depth depending on the data (as reported below) and using batch normalization and ReLU activation after each layer except for the output layer.
@@
For $\hat{h}$, we set $X$ to be the latent vectors of the training set (the encoded images), and set $Y$ to be the standard multivariate normal distribution of the same dimension as the latent space (e.g., in CelebA we used a latent space of $100$ in the autoencoder, so $Y$ has dimension $100$). Images were generated using Algorithm 2 by computing: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ where $r\sim\mathcal{N}(0,I)$ (generation occurs in the latent space).
@@
Figure 4 shows random sets of generated images for MNIST at several epochs of improvement on the validation set, as well as a sample of random real images for comparison. These results show that the generated samples were already realistic even after one epoch (a similar observation was also made in the swiss-roll example). We use a latent dimension of $5$ to allow for comparison with VAE for the same latent dimension (Figure 5(b) in (Kingma and Welling 2013)). Further comparisons can be made with diffusion (Figure App.1. in (Sohl-Dickstein et al. 2015b)) and a recent continuous normal flow (CNF) (Figure 3(a) (right) in (Huang and Yeh 2021)). For $\hat{h}$ we used width $50$ and depth $5$. Training was stopped when no improvement in the validation MSE was obtained for $3$ epochs.
@@
Next we applied the method to CelebA (Figure 1 and Appendix Figure 7). Appendix Figure 7 also shows the progress in image generation during training, with randomly generated images at epochs of improvement in the Inception Score (IS) (Salimans et al. 2016). We used IS as the stopping criteria, as explained in Appendix A.4. One epoch took just under $1$ minute ($50.6$ seconds on average) on a single T4 GPU, reflecting $9$ hours until the last improvement in IS at epoch $640$. Observing Appendix Figure 7, we can see that realistic images were already obtained at half this time (see epoch $276$). Appendix Figure 8 contains a plot of the IS by epoch showing that the true IS (the IS score for a random set of real decoded samples) is nearly achieved by the GTN. The time to generate $1,000$ samples (i.e. to compute: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$) on a standard PC was $4.07$ seconds (reflecting $4.07$ milliseconds per image) and $0.00367$ seconds (reflecting $0.00367$ milliseconds per image) on a T4 GPU. We used a latent dimension of $100$ in the autoencoder. For $\hat{h}$ we used width $1200$ and depth $25$. We used a center crop of $148\times 148$ and resized the images to $64\times 64$.
Interpolation.
Figure 5: Interpolations produced by a trained GTN for CelebA (top) and the HaP dataset (bottom).
@@
We next used the CelebA dataset as well as the Hands and Palm Images (HaP) dataset (Kaggle. 2024) to demonstrate that GTNs indeed provide continuous interpolations. The HaP dataset is a small dataset containing $11,000$ images of upward-facing and downward-facing hands. We chose the HaP dataset since generating hands is known to be notoriously difficult for generative models (The New Yorker. 2023; Vox. 2023), let alone interpolating between them.
@@
Figure 5 shows that GTN indeed generates continuous interpolations for both CelebA and HaP. Each grid contains generated interpolations between two generated images (top left and bottom right in each grid). More precisely, each grid contains the results of decoding $\hat{h}\big{(}\lambda y_{\text{top\_left}}+(1-\lambda)y_{\text{bottom\_right}}\big{)}$ for $100$ linearly spaced $\lambda\in[0,1]$ where both $y_{\text{top\_left}}$ and $y_{\text{bottom\_right}}$ were sampled from $\mathcal{N}(0,I)$.
@@
It is worth noting an interesting observation for the HaP data results – the interpolations in Figure 5 are between hands of the same orientation (downward facing). However, generated interpolations between hands (downward facing) and palms (upward-facing) do not seem as natural (Appendix Figure 9). This is likely because the dataset does not contain in-between positions for the latter (transitions between downward and upward), but it does contain in-between positions for the former (closed/open fingers to various degrees). This demonstrates that, regardless of methodology, the completeness of the dataset is important for accurate image generation, and particularly for interpolation.
@@ 4 Related work
GTNs are related in concept to NFs which seek to map a source distribution into a target distribution using a sequence of bijective transformations. These are typically implemented by a neural network, often in the form of at least tens of neural network blocks (Xu and Campbell 2024) and sometimes more (Xu, Chen, and Campbell 2023). The core of NFs is based on each of these transformations being a diffeomorphism – a notion similar to a homeomorphism but more constrained as it is defined between smooth manifolds (as opposed to any topological spaces, including any manifolds) and requires that the function and its inverse are differentiable (as opposed to just continuous for homeomorphisms). NFs also require that the log-determinant of the Jacobian of these transformations is tractable, posing a constraint on the model architecture. GTNs do not pose any limitations on the model architecture. NFs also differ from GTNs in their optimization method since NFs employ a maximum likelihood objective – typically the Kullback-Leibler (KL) divergence while GTNs are trained using MSE. It was observed that optimizing the KL-divergence may be difficult for more complex normalizing flow structures (Xu, Chen, and Campbell 2023). NFs may also suffer from low-quality sample generation (Behrmann et al. 2021), partially because the constraint on the Jacobian, besides limiting the model structure, may also lead to issues such as spurious local minima and numerical instability (Golinski, Lezcano-Casado, and Rainforth 2019).
Continuous normalizing flow (CNF) are a relatively recent type of NFs that were developed to avoid the main constraints posed by NFs – namely the requirements of invertibility and a tractable Jacobian determinant. To avoid these constraints, CNFs use ODEs to describe the transformations in NFs. While CNFs have shown improvements over NFs in certain cases , they are extremely time-intensive since they require solving a series of ODEs, equivalent to running a neural network with hundreds of layers (Huang and Yeh 2021). Improving both the speed and performance of CNFs is an active field of research. To compare between a recently suggested CNF and GTN see the Experiments section.
@@ 5 Discussion
In this work we introduced a new class of generative models – GTNs – that learn a homeomorphism $h$ between the supports of two continuous random variables using samples from both. The samples are used in an efficient algorithm to create a training dataset for the GTN which is trained in a simple supervised approach, with no restrictions on the model architecture. The freedom to use a standard neural network architecture, and the fact that $h$ satisfies the conditions of the universal approximation theorem, lead GTNs to quickly converge towards $h$, providing an efficient method to generate samples from a target random variable. This was demonstrated experimentally in multiple cases, particularly in that generated samples reached the data manifold at early epochs and in the total time to convergence. Other positive implications of $h$ being a homeomorphism – namely, coverage and diversity, continuous interpolations and guiding topological properties – were also discussed and demonstrated experimentally.
One downside of our method is that it assumes that the data is in its intrinsic dimension (or lower). This may require use of dimensionality reduction methods, such as the vanilla autoencoders employed in this work. Future work focused on high resolution images may be interested in experimenting with stronger autoencoders. Additionally, the assumptions on the support of the distributions mean that our method does not immediately apply to discrete distributions, and that datasets may need to be split and trained on individually, as explained in Appendix A.3. This also means that larger datasets may be required for GTNs to perform well.
6 Broader Impact
Our work falls under the category of deep generative models, and is similar in impact to other studies seeking to improve these methods. Since our work provides an efficient method to generate realistic samples, it may contribute to malicious uses such as generating fake news, videos of high profile figures etc. Fortunately, there are methods to detect such generated data, and there is ongoing work to continue to improve in doing so (Wang et al. 2020). There are also concerns of perpetuating biases inherent to the dataset, particularly since our method essentially learns the data distribution with high fidelity. There is a growing line of research on detecting and handling biases in datasets, and we believe there is an increasing awareness among data curators and users of this problem. More research would be necessary in order to alleviate this problem.
Although malicious uses are indeed a concern, generative models could have far reaching positive implications on our lives. Perhaps the most prominent example is the field of healthcare, where generative models have found use in generating digital twins for improved personalized medicine (Bordukova et al. 2024; Tsialiamanis et al. 2021), medical imaging (Kazerouni et al. 2023; Moghadam et al. 2023; Pinaya et al. 2022) and drug design (Korablyov et al. 2024; Chen et al. 2023; Hoogeboom et al. 2022; Schneuing et al. 2022; Corso et al. 2022).
Appendix A Appendix
A.1
@@ Appendix A Appendix
    $h$ is well-defined: we need to show that: (a) ${F_{X}}|_{S_{X}}$ is invertible, and (b) that the image of ${F_{Y}}|_{S_{Y}}$ is in the domain of ${F_{X}}|_{S_{X}}$. For (a), observe that ${F_{X}}|_{S_{X}}$ is continuous, strictly monotonically increasing (the CDF is strictly monotonically increasing on the support) and onto $(0,1)$. A symmetric argument can be made for ${F_{Y}}|_{S_{Y}}$, which we will use later to define the inverse of $h$.
    For (b), we know from (a) that the image of ${F_{Y}}|_{S_{Y}}$ is the domain of ${F_{X}}|_{S_{X}}$ so $h$ is well defined.
@@
    $h$ is a homeomorphism since: a. it is continuous as a composition of continuous functions (${F_{X}^{-1}}|_{S_{X}}$ is the inverse of a continuous and strictly monotonous function on an open interval in $\mathbb{R}$); b. it is bijective with inverse ${F_{Y}^{-1}}|_{S_{Y}}({F_{X}}|_{S_{X}}(x))$; c. its inverse is continuous for the same reasons stated in a. Also, it is a known result that a continuous and bijective function between two open intervals has a continuous inverse.
A.2
To understand the rationale behind Algorithm 1, imagine four points from each of $X$ and $Y$ lying on the $x$-axis in $\mathbb{R}^{2}$ such that both $X$ and $Y$ have two points on each side of the origin (we assume $D_{X},D_{Y}$ are centralized). The cosine similarity of each pair of points $\{a,b\}$ with $a\in D_{X},b\in D_{Y}$ is $1$ if they are on the same side or $-1$ of they are on opposite sides. Take the first $y\in D_{Y}^{\text{sorted}}$ (the closest one to the origin which reflects $\sim 50$th percentile since we assumed an equal number of points from $Y$ on both sides). The maximum cosine similarity out of all $x\in D_{X}^{\text{sorted}}$ is $1$, meaning $y$ will be matched with an $x$ on the same side. The fact that $D_{X}$ is sorted, means that we are using the distance as a tie breaker – out of all $x\in D_{X}$ that are on the same side as $y$, the $x$ that is closest to the origin will be chosen ($\sim 50$th percentile). Likewise, the leftmost $y$ will be matched with the leftmost $x$ (both 00th percentile), the rightmost $y$ with the rightmost $x$ (both $100$th percentile) etc.
A.3
One example of a topological property is the connectedness property – being a ”single piece”, like $\mathbb{R}$, as opposed to ”more than one piece”, like $(-\infty,-1)\cup(1,\infty)$, is preserved between homeomorphic spaces (so, in particular, these two are not homeomorphic, but $\mathbb{R}$ is homeomorphic to each of the two intervals, separately). Therefore, if we suspect that our data is composed of separate classes like images of both hands and faces, and we wish $\hat{h}$ as a generative model, we could split our data to its different components (learning a separate $h$ for each). This relates to the conditions of Theorem 2.1 which assume that $X$’s support is an open interval. For example, if we suspect that our data can be separated into two disjoint uniform distributions, then the theorem doesn’t apply. However, it does apply to each separate uniformly distributed component.
@@
Figure 7: Randomly chosen samples generated by GTN during the training process on celebA at several epochs of improvement in the IS score, and randomly chosen real images (bottom-right grid). Each epoch shows 200 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(0,I)$). The bottom right grid contains 200 images each of which is the decoded latent vector of a random real image.
A.4
We experimented with two stopping criteria. The first was based on the validation set, as in the MNIST case. The second used IS to assess the quality of the generated images using a small random set of $200$ images. IS yielded better results. Training was stopped when no improvement in the IS score was obtained for $300$ epochs, but observing the generated samples by epoch above demonstrates that earlier stopping may suffice.
Figure 8: Inception score (IS) for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors) plotted by epoch (at epochs of improvement in the score) (blue); and IS of a random set of $200$ real images (decoded latent vectors of real images) (green).
Figure 9: Interpolations between a hand (downward facing) and a palm (upward facing) (left) and vice versa (right).
A.5
$h$ is injective as can be observed by taking the norm when equating for two different points $y_{1},y_{2}$:
@@
which yields $h_{1}(||y_{1}||)=h_{1}(||y_{2}||)$. Since $h_{1}$ is injective, this means that $||y_{1}||=||y_{2}||$. Plugging this into the equality and cancelling out equal terms ($y_{1},y_{2}\neq 0)$ yields $y_{1}=y_{2}$. The function is also onto since $h_{1}$ is onto. Particularly, for any $x\neq 0$ there is a value $v$ in $(0,\infty)$ with $h_{1}(v)=||x||$. Since the support of $Y$ is $\mathbb{R}^{d}$ there is a $y$ that is on the same line as $x$ from the origin and that satisfies $||y||=v$ so that: $x=h_{1}(||y||)\frac{y}{||y||}$.
$h$ is continuous since for $y\neq 0$ it is a composition of continuous functions, and for $y=0$ we can observe that $\lim_{y\to 0}h_{1}(||y||)\frac{y}{||y||}=0$ ($\frac{y}{||y||}$ is the unit vector in the direction of $y$ and $h_{1}(||y||)\to 0$ as $y\to 0$ by definition of the CDF).

### v2

@@ 3 Experiments
In light of the discussion in Section 2.1 on using the intrinsic dimension (ID) of the data, we use autoencoders to represent each dataset in a lower dimensional latent space before training $\hat{h}$. Each dataset was trained separately (with its own autoencoder and $\hat{h}$). In all experiments, we use the same vanilla autoencoder architecture adapted to different latent dimensions (the encoder consisted of $2$ convolution layers with ReLU activation followed by a fully connected layer). For $\hat{h}$ we use a standard feedforward neural network with the width and depth depending on the data. All architecture and training details for all experiments are in Appendix F.
@@
To train $\hat{h}$, we set $X$ to be the latent vectors of the training set (the encoded images), and set $Y$ to be the standard multivariate normal distribution of the same dimension as the latent space (e.g., if we used a latent dimension of $d=100$ in the autoencoder, then $Y$ has dimension $100$). Images were generated using Algorithm 2 by computing: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ where $r\sim\mathcal{N}(\textbf{0},\textbf{I})$ (generation occurs in the latent space).
Results.
@@
We first apply our method to MNIST (LeCun et al. 1998) using a latent dimension of $5$. We chose $d=5$ to allow for comparison with VAE for the same latent dimension (Figure 5(b) in (Kingma and Welling 2013)). Figure 5 shows random sets of generated images for MNIST at several epochs of improvement on the validation set, as well as a sample of random real images for comparison. These results show that the generated samples were already realistic even after one epoch (a similar observation was made for the swiss-roll). Besides the aforementioned comparison to VAE, further comparisons can be made with diffusion (Figure App.1. in (Sohl-Dickstein et al. 2015b)) and a continuous normal flow (CNF) (Figure 3(a) (right) in (Huang and Yeh 2021)). Training was stopped when no improvement in the validation MSE was obtained for $3$ epochs.
Next we applied the method to CelebA (Liu et al. 2015) (Figure 1 and Appendix Figures 10-14). Images were center cropped to $148\times 148$ and resized to $64\times 64$ We tested latent dimensions of $d\in\{10,25,50,100\}$, which are within the range of the typical ID estimated for image datasets ($10$-$50$) or close ($100$) (Pope et al. 2021). We used Inception Score (IS) (Salimans et al. 2016), a common evaluation metric, as the stopping criteria: after every epoch, the GTN generated $200$ random images for which the IS was evaluated. We used IS instead of the validation set since the IS continued to improve well beyond the point of plateau on the validation set (for MNIST the IS was irrelevant since the data is grayscale). Training stopped after $300$ epochs of no improvement in IS. We also produced the IS for a single random set of $200$ real decoded images as a baseline (”real IS”). Plotting the IS by epoch, alongside the real IS (Appendix Figure 9) shows that the real IS is either achieved ($d=10,25$) or nearly achieved ($d=50,100$) by the GTN and that the gap between IS and real IS decreases as $d$ decreases. With the final model, we also computed another common evaluation metric – the Fréchet Inception Distance (FID) (Heusel et al. 2017) – between the training set and an equally-sized set of generated images. We observed a similar pattern, namely that FID decreased as $d$ decreased, except for $d=10$) (Appendix Table 2). Overall, the IS and FID, as well as the difference between IS and real IS, improved as $d$ decreased. Visual inspection of randomly generated images next to real images across all dimensions (Appendix Figures 10-13) confirms these improvements. Namely, as $d$ decreases we see more similarity between generated images and real images as well as improved diversity.
@@
Observing the progress in image generation during training (Appendix Figure 14), we can see that realistic images were already obtained at half the training time (see epoch $276$). One epoch took just under $1$ minute ($50.6$ seconds on average) on a single T4 GPU, reflecting $9$ hours until the last improvement in IS at epoch $640$ and $<4$ hours to reach epoch $276$. The time to generate $1,000$ samples (i.e. to compute: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$) on a standard PC was $4.07$ seconds (reflecting $4.07$ milliseconds per image) and $0.00367$ seconds (reflecting $0.00367$ milliseconds per image) on a T4 GPU.
Figure 6: Interpolations generated by a trained GTN for CelebA (top) and the HaP dataset (bottom).
@@
We next used the CelebA dataset as well as the Hands and Palm Images (HaP) dataset (Kaggle. 2024) to demonstrate that GTNs indeed provide continuous interpolations. The HaP dataset is a small dataset containing $11,000$ images of upward-facing and downward-facing hands. We chose the HaP dataset since generating hands is known to be notoriously difficult for generative models (The New Yorker. 2023; Vox. 2023), let alone interpolating between them. As in CelebA, we trained a GTN for each of the four dimensions, on images resized to $64\times 64$.
Figure 6 shows interpolations for both CelebA and HaP. For CelebA we used the $d=100$ model and for HaP we used $d=50$. The dimensions were chosen after consulting the IS and FID results across all four $d$ values in Appendix Table 2 while preferring a higher dimension ($d\in\{50,100\}$) to retain more visual detail (e.g., compare Appendix Figures 10-11 to 12-13): in CelebA $d=100$ obtained similar results to $d=50$ and in HaP $d=50$ obtained better results than $d=100$. Note that we observe similar trends in HaP as in CelebA: the difference between IS and real IS, as well as the FID decrease with $d$ (Appendix Figure 16 and Table 2).
@@
Figure 6 shows that GTN indeed generates continuous interpolations for both CelebA and HaP. Each row contains generated interpolations between two images (leftmost and rightmost). Specifically, each row contains the results of decoding $\hat{h}\big{(}\lambda y_{\text{left}}+(1-\lambda)y_{\text{right}}\big{)}$ for $20$ linearly spaced $\lambda\in[0,1]$, where, in CelebA, both $y_{\text{left}}$ and $y_{\text{right}}$ were sampled from $\mathcal{N}(\textbf{0},\textbf{I})$, and in HaP $y_{\text{left}}$ and $y_{\text{right}}$ are each the normal label of a real encoded image (from Algorithm 1), chosen so that they have the same orientation.
@@
It is worth noting an interesting observation arising from the HaP dataset. Notice that the generated interpolations in Figure 6 are between hands of the same orientation (downward facing). However, generated interpolations between hands (downward facing) and palms (upward-facing) do not seem as natural (Appendix Figure 15). This is likely because the dataset does not contain in-between positions for the latter (transitions between downward and upward), but it does for the former (closed/open fingers to various degrees). This demonstrates that, regardless of methodology, the completeness of the dataset is important for accurate image generation, and particularly for interpolation.
@@ 4 Related work
GTNs are related in concept to NFs which seek to map a source distribution into a target distribution using a sequence of bijective transformations. These are typically implemented by a neural network, often in the form of at least tens of neural network blocks (Xu and Campbell 2024) and sometimes more (Xu, Chen, and Campbell 2023). The core of NFs is based on each of these transformations being a diffeomorphism – a specific type of homeomorphism that is more constrained as it is defined between smooth manifolds (as opposed to any topological spaces, including any manifolds) and requires that the function and its inverse are differentiable (as opposed to just continuous for general homeomorphisms). NFs also require that the log-determinant of the Jacobian of these transformations is tractable, posing a constraint on the model architecture. GTNs do not pose any limitations on the model architecture. NFs also differ from GTNs in their optimization method since NFs employ a maximum likelihood objective – typically the Kullback-Leibler (KL) divergence while GTNs are trained using MSE. It was observed that optimizing the KL-divergence may be difficult for more complex normalizing flow structures (Xu, Chen, and Campbell 2023). NFs may also suffer from low-quality sample generation (Behrmann et al. 2021), partially because the constraint on the Jacobian, besides limiting the model structure, may also lead to issues such as spurious local minima and numerical instability (Golinski, Lezcano-Casado, and Rainforth 2019).
Continuous normalizing flow (CNF) are a relatively recent type of NFs that were developed to avoid the main constraints posed by NFs – namely the requirements of invertibility and a tractable Jacobian determinant. To avoid these constraints, CNFs use ordinary differential equations (ODEs) to describe the transformations in NFs. While CNFs have provided certain improvements over NFs, they are still time-intensive since they require solving a series of ODEs, equivalent to running a neural network with hundreds of layers (Huang and Yeh 2021). Improving both the speed and performance of CNFs is an active and promising field of research. A comparison between a recent CNF and GTN was suggested in the Experiments section.
@@ 5 Discussion
In this work we introduced a new class of generative models – GTNs – that learn a homeomorphism $h$ between the supports of two continuous random variables using samples from both. The samples are used in an efficient algorithm to create a training dataset for the GTN which is trained in a simple supervised approach, with no restrictions on the model architecture. The freedom to use a standard neural network architecture, and the fact that $h$ satisfies the conditions of the universal approximation theorem, lead GTNs to quickly converge towards $h$, providing an efficient method to generate samples from a target random variable. This was demonstrated experimentally in multiple cases, particularly manifesting in the observation that generated samples reached the data manifold at early epochs and in the total time to convergence. Other positive implications of $h$ being a homeomorphism – namely, coverage and diversity, continuous interpolations and guiding topological properties – were also discussed and demonstrated experimentally.
One downside of our approach is that it relies on the theoretical assumption that the data is represented in its intrinsic dimension. Experimentally, this assumption manifested in different performances and sample quality across different dimensions. This means that dimensionality reduction methods, like the autoencoders used in this work, may be essential for high quality sample generation. Future work, e.g. for generating high-resolution images, may be interested in experimenting with stronger autoencoders as well as additional procedures such as data augmentations, which were not used here. Additionally, the assumptions on the supports of the distributions mean that our method may be ill-suited for discrete distributions, and that datasets may need to be split into separate components for individual training, as explained in Appendix C. The latter case means that larger datasets may be required to ensure that the number of samples per component is sufficiently large. Existing literature on estimating the intrinsic dimension of the data and on the relation between the intrinsic dimension and the required sample complexity (e.g., (Pope et al. 2021)) may be of value in both of these contexts.
Appendix
Appendix A
@@
    $h$ is well-defined: it suffices to show that: (a) $F_{X}|_{S_{X}}$ is invertible, and (b) that the image of $F_{Y}|_{S_{Y}}$ is in the domain of $F_{X}|_{S_{X}}^{-1}$. For (a), observe that $F_{X}|_{S_{X}}$ is continuous (the CDF of a continuous random variable is continuous) , strictly monotonically increasing (the CDF is strictly monotonically increasing on the support) and onto $(0,1)$ (by definition of the CDF and the fact that the support is an open interval). As such it is invertible. A symmetric argument can be made for $F_{Y}|_{S_{Y}}$, which we will use later to define the inverse of $h$.
    For (b), we know from the proof of (a) that the image of $F_{Y}|_{S_{Y}}$ is the domain of $F_{X}|_{S_{X}}^{-1}$.
@@
    $h$ is a homeomorphism since: a. it is continuous as a composition of continuous functions. Indeed, $F_{Y}|_{S_{Y}}$ is continuous (the CDF of a continuous random variable is continuous). $F_{X}|_{S_{X}}^{-1}$ is also continuous: since $F_{X}|_{S_{X}}$ is continuous (as was $F_{Y}|_{S_{Y}}$) and bijective (it is invertible, as shown in the first part) we can use the known result that a continuous and bijective function between two open intervals has a continuous inverse (a consequence of the invariance of domain theorem). ; b. it is bijective since the inverse is $F_{Y}|_{S_{Y}}^{-1}\circ F_{X}|_{S_{X}}$; c. its inverse is continuous: we know from a. and b. that $h$ is continuous and bijective so we can again use the fact that a continuous and bijective function between two open intervals has a continuous inverse.
Appendix B
To understand the rationale behind Algorithm 1, imagine four points from each of $X$ and $Y$ lying on the same line, say the $x$-axis in $\mathbb{R}^{2}$, such that both $X$ and $Y$ have two points on each side of the origin (we assume $D_{X},D_{Y}$ are centralized). See Appendix Figure 7, where circles are from $Y$ and stars are from $X$. The cosine similarity of each pair of points $\{a,b\}$ with $a\in D_{X},b\in D_{Y}$ is $1$ if they are on the same side or $-1$ of they are on opposite sides. Take the first $y\in D_{Y}^{\text{sorted}}$ (the closest one to the origin which reflects $\sim 50$th percentile since we assumed an equal number of points from $Y$ on both sides). The maximum cosine similarity out of all $x\in D_{X}^{\text{sorted}}$ is $1$, meaning $y$ will be matched with an $x$ on the same side. The fact that $D_{X}$ is sorted, means that we are using the distance as a tie breaker – out of all $x\in D_{X}$ that are on the same side as $y$, the $x$ that is closest to the origin ($\sim 50$th percentile) will be chosen as $x_{y}$. Likewise, the leftmost $y$ will be matched with the leftmost $x$ (both 00th percentile), the rightmost $y$ with the rightmost $x$ (both $100$th percentile) etc. Note that if this were in $\mathbb{R}$, this process coincides with the same labeling process described in the $1$D case (illustrated in Figure 2).
Figure 7: Illustration of Algorithm 1 as described in Appendix B. The numbers reflect the order in which the matching of $y$ with $x_{y}$ occurs.
Appendix C
One example of a topological property is the connectedness property – being a ”single piece”, like $\mathbb{R}$, as opposed to ”more than one piece”, like $(-\infty,-1)\cup(1,\infty)$, is preserved between homeomorphic spaces (so, in particular, these two are not homeomorphic, but $\mathbb{R}$ is homeomorphic to each of the two intervals, separately). Therefore, if we suspect that our data is composed of separate classes like images of both hands and faces, and we would like to train a GTN as a generative model, we could split our data to its different components (learning a separate $h$ for each). This relates to the conditions of Theorem 2.1 which assume that $X$’s support is an open interval. For example, if we suspect that our data can be separated into two disjoint uniform distributions, then the theorem doesn’t apply. However, it does apply to each separate uniformly distributed component.
@@
Figure 9: Inception score (IS) vs real IS for CelebA plotted by epoch (at epochs of improvement in the IS). At each epoch, the IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors, formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). Real IS was computed once for a random set of $200$ real images (decoded latent vectors of real images) (dashed gray line).
Figure 10: Generated vs real samples for latent dimension $d=10$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 11: Generated vs real samples for latent dimension $d=25$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 12: Generated vs real samples for latent dimension $d=50$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 13: Generated vs real samples for latent dimension $d=100$. (A) Random generated samples – each image is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). (B) Random real decoded images – each image is the decoded latent vector of a random real image.
Figure 14: Randomly chosen samples generated by a GTN during the training process on CelebA with latent dimension 100 at several epochs of improvement in the IS score, and randomly chosen real images (bottom-right grid). Specifically, each epoch shows 200 images, each of which is the decoded $\hat{h}(r)$ for some random $r\sim\mathcal{N}(\textbf{0},\textbf{I})$). The bottom right grid shows 200 images each of which is the decoded latent vector of a random real image. Training until epoch 640 took approximately $9$ hours on a single T4 GPU, with similar-quality images obtained in less than $4$ hours of training at epoch $276$.
Figure 15: Interpolations between a hand (downward facing) and a palm (upward facing).
Figure 16: Inception score (IS) vs real IS for HaP plotted by epoch (at epochs of improvement in IS). At each epoch, IS was computed for $200$ randomly generated images (decoded latent vectors generated by $\hat{h}$ from random normal vectors), formally: $autoencoder.decoder\big{(}\hat{h}(r)\big{)}$ for $r\sim\mathcal{N}(\textbf{0},\textbf{I})$) (solid black line). Real IS was computed once for a random set of $200$ real images (decoded latent vectors of real images) (dashed gray line).
Appendix D
$h$ is injective: equating for two points $y_{1},y_{2}$ we obtain:
@@
Applying the norm to both sides yields $h_{1}(||y_{1}||)=h_{1}(||y_{2}||)$. Since $h_{1}$ is injective, this means that $||y_{1}||=||y_{2}||$. Plugging this into the equality and cancelling out equal terms ($y_{1},y_{2}\neq 0)$ yields $y_{1}=y_{2}$. The function is also onto: since $h_{1}$ is onto (it is a homeomorphism) then for any $x\neq 0$ there is a value $v$ in $(0,\infty)$ with $h_{1}(v)=||x||$. Since the support of $Y$ is $\mathbb{R}^{d}$ there is a $y$ that is on the same line as $x$ from the origin and that satisfies $||y||=v$ so that: $x=h_{1}(||y||)\frac{y}{||y||}$, meaning that $x$ has a $y\in S_{Y}$ for which $h(y)=x$.
$h$ is continuous since for $y\neq 0$ it is a composition of continuous functions, and for $y=0$ we observe that $\lim_{y\to 0}h_{1}(||y||)\frac{y}{||y||}=0$ ($\frac{y}{||y||}$ is the unit vector in the direction of $y$ and $h_{1}(||y||)\to 0$ as $y\to 0$ by definition of the CDF).